OpenStudy (anonymous):

Use complete sentences to describe a real-world scenario that could be represented by the inequality 4x + 2y 40.

OpenStudy (anonymous):

4x+2y ???40

OpenStudy (anonymous):

what inequality

OpenStudy (radar):

\[\ge \le ><\neq ????\]

OpenStudy (anonymous):

greater then or equal to

OpenStudy (anonymous):

\[\ge\]

OpenStudy (anonymous):

your illegal bar sells whiskey at $4 a shot, beer at $2 a glass. how much do you have to sell to collect more than $40?

OpenStudy (anonymous):

if you don't think your teacher would appreciate that, change whiskey to pocky and beer to hershey bars

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

o.0 thanks :P